OpenStudy (anonymous):

What is the common difference in the arithmetic sequence 30, 27, 24, 21, 18, ...?

OpenStudy (anonymous):

second term - first term = 27-30 = -3

OpenStudy (anonymous):

or 3rd term - 2nd term = 24-27 = -3

OpenStudy (anonymous):

or the n+1st term - nth term